The table below gives the age and bone density for five randomly selected women. Using this data, consider the equation of the regression line, yˆ=b0+b1xy^=b0+b1x, for predicting a woman's bone density based on her age. Keep in mind, the correlation coefficient may or may not be statistically significant for the data given. Remember, in practice, it would not be appropriate to use the regression line to make a prediction if the correlation coefficient is not statistically significant.

Age 4242 4343 5151 5656 6565
Bone Density 353353 344344 333333 324324 322322

Table

Copy Data

Step 6 of 6 :  

Find the value of the coefficient of determination. Round your answer to three decimal places.
Answer #1

Step 6:

X Values
∑ = 257
Mean = 51.4
∑(X - Mx)2 = SSx = 365.2

Y Values
∑ = 1676
Mean = 335.2
∑(Y - My)2 = SSy = 698.8

X and Y Combined
N = 5
∑(X - Mx)(Y - My) = -471.4

R Calculation
r = ∑((X - My)(Y - Mx)) / √((SSx)(SSy))

r = -471.4 / √((365.2)(698.8)) = -0.933

So coefficient of determination is r^2=-0.933^2=0.870